OpenStudy (anonymous):

PLEASE HELP! Show all work in dividing (y^2-3y+2)/(y^2+5y+6) by (y^2+2y-8)/(y^2+y-12) . Use complete sentences to explain how to simplify this expression.

OpenStudy (anonymous):

factorize the equations in numerator and denominator, Y^2-3y+2=(y-1)(y-2) and y^2+5y+6=(y+2)(y+3) do the same for the next fraction and then multiply and cancel out the factors. :)